### v1

# Separable commutative algebras in equivariant homotopy theory

Abstract.
Given a finite group $G$ and a commutative ring $G$-spectrum $R$, we study the separable commutative algebras in the category of compact $R$-modules. We isolate three conditions on the geometric fixed points of $R$ which ensure that every separable commutative algebra is standard, i.e. arises from a finite $G$-set. In particular we show that all separable commutative algebras in the categories of compact objects in $G$-spectra and in derived $G$-Mackey functors are standard provided that $G$ is a $p$-group. In these categories we also show that for a general finite group $G$, we cannot expect all separable commutative algebras to be standard.
2020 Mathematics Subject Classification:
18F99, 55P42, 55P91, 55U35
1. Introduction
The study of separable commutative algebras in the setting of tensor-triangular geometry was initiated by Balmer in [Bal11]. Separable commutative algebras are useful in this context for several reasons. Firstly, they allow to form categories of modules internally to the world of triangulated categories, without the necessity of assuming that the triangulated category admits a model. Building on this, Balmer showed that restriction functors in (modular) representation theory can be understood as an extension-of-scalars functors [Bal15], and proved a generalization of the Neeman-Thomason Localization Theorem [Bal16a]. Secondly, separable commutative algebras allow us to formulate a going-up theorem in tt-geometry [Bal16b], which is a useful tool for calculating the Balmer spectrum of a triangulated category, and powerful descent results in tt-geometry [Bal16b],[BCH+24] and [BCH+23]. Finally, separable commutative algebras are tightly related to the étale topology of the given tt-category. A result of Neeman [Nee18] informally says that the only separable commutative algebras in the perfect derived category of a Noetherian scheme are those arising from finite étale morphisms. This connection with the étale topology was strengthened even further by the first two authors in [NP24] where they showed that separable commutative algebras recover finite étale algebras in various contexts where a notion of étaleness already exists.
The study and classification of separable commutative algebras is more complicated when some equivariance comes into play. Given a finite group $G$ and a separably closed field $k$, Balmer showed that the only separable commutative algebras in the category of $kG$-modules are those of the form $kX$ for some finite $G$-set $X$. We will informally refer to this result as saying that in the category of $kG$-modules all separable algebra are standard. Balmer went one step further and asked if the same holds in the stable module category [Bal15, Question 4.7]:
Question (Balmer).
Let $k$ be a separably closed field and $A\in\operatorname{stmod}(kG)$ be a separable commutative algebra. Is there a finite $G$-set $X$ such that $A\simeq kX$ in $\operatorname{stmod}(kG)$?
Informally this problem ask whether the étale topology which appears in modular representation theory is richer than what is produced by subgroups. A positive answer to this question was given by Balmer-Carlson [BC18] in the case of cyclic $p$-groups and by the first two authors [NP24] for $p$-groups of $p$-rank one. This question is completely open outside of these cases.
In this paper we aim to address the analogue of Balmer’s question in the setting of equivariant stable homotopy theory. Fix a finite group $G$ and a commutative ring $G$-spectrum $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$, i.e. a commutative algebra in genuine $G$-spectra. Given a finite $G$-set $X$, we can form the commutative algebra $\mathbb{D}(X_{+})\in\operatorname{CAlg}(\operatorname{Sp}_{G})$ by taking the dual of the suspension spectrum of $X_{+}$; this is separable by work of Balmer-Dell’Ambrogio-Sanders [BDS15]. Base-changing to our $R$ we obtain a functor
$R^{(-)}\colon\operatorname{Fin}^{\operatorname{op}}_{G}\to\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega})$
into the category of separable commutative algebras in compact $R$-modules. We will denote the restriction to groupoids of this functor as
$R^{(-)}\colon\mathcal{F}\mathrm{in}^{\operatorname{op}}_{G}\to\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega}).$
In this paper we isolate three conditions that ensure that this functor is an equivalence, showing in particular that every separable algebra is standard. Recall that for all subgroups $K\subseteq G$, the $K$-geometric fixed points $\Phi^{K}(R)$ carries an action of the Weyl group $W_{G}(K)$ so that
$\Phi^{K}(R)\in\operatorname{Fun}(BW_{G}(K),\operatorname{CAlg}).$
We then consider the following conditions for $(R,K)$:
  • (1)
    the commutative algebras $(\Phi^{K}(R))^{hW_{G}(K)}$ and $(\Phi^{K}(R))^{tW_{G}(K)}$ are indecomposable.
  • (2)
    If $\Phi^{K}(R)$ is a retract in $\operatorname{Mod}(\Phi^{K}(R))^{tW_{G}(K)}$ of the coinduced algebra $(\Phi^{K}(R))^{W_{G}(K)/U}$ for some $U\subseteq W_{G}(K)$, then $U=W_{G}(K)$, see 6.1.
  • (3)
    $\Phi^{K}(R)$ is separably closed, see 7.1.
We will show that the above conditions are satisfied if $R=\mathbb{S}_{G}$ is the sphere $G$-spectrum, and also if $R=\operatorname{infl}_{e}^{G}\mathbb{Z}$, provided that $G$ is a nontrivial $p$-group. The latter case has gained increasing interest for its connection to the category of derived $G$-Mackey functors of Kaledin, see [PSW22]. Our main result then states:
Theorem (9.6).
Suppose that $G$ and $R$ satisfy conditions (1), (2) and (3) for all $K\subseteq G$. Then
$R^{(-)}\colon\mathcal{F}\mathrm{in}^{\operatorname{op}}_{G}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega})$
is an equivalence.
It follows that if $G$ is a $p$-group, then all separable commutative algebras are standard, both in compact $G$-spectra and in compact derived $G$-Mackey functors. As an immediate application, we also show that the Galois groups of these categories, in the sense of [Mat16], are trivial, see 9.8. The techniques developed in this paper also allow us to produce some counterexamples. Indeed we show in Section 10 that if $G=C_{6}$, then there are separable commutative algebras in the above two categories which are not standard.
Acknowledgements
We thank Bastiaan Cnossen, John Greenlees and Achim Krause for useful conversations.
The first and second authors are supported by the SFB 1085 Higher Invariants in Regensburg. The third author was supported by the Danish National Research Foundation through the Copenhagen Centre for Geometry and Topology (DNRF151) by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Project-ID 427320536 - SFB 1442, as well as under Germany’s Excellence Strategy EXC 2044 390685587, Mathematics Münster: Dynamics-Geometry-Structure. .
Notations
  1. (a)
    We denote by $\operatorname{Cat}_{\infty}$ the $\infty$-category of $\infty$-categories.
  2. (b)
    We denote by $\operatorname{CAlg}$ the $\infty$-category of $\mathbb{E}_{\infty}$-algebras in spectra and by $\operatorname{CAlg}^{\heartsuit}$ the category of discrete commutative rings.
  3. (c)
    We denote by $\operatorname{Cat}_{\infty}^{\mathrm{st}}$ the $\infty$-category of essentially small stable $\infty$-category and exact functors.
  4. (d)
    We denote by $\operatorname{CAlg}(\operatorname{Cat}_{\infty}^{\mathrm{st}})$ the $\infty$-category of essentially small, symmetric monoidal stable $\infty$-categories whose tensor product is exact in each variable, and symmetric monoidal exact functors.
  5. (e)
    We denote by $\operatorname{Cat}_{\infty}^{\mathrm{perf}}$ the $\infty$-category of essentially small, idempotent complete stable $\infty$-categories and exact functors.
  6. (f)
    We denote by $\operatorname{CAlg}(\operatorname{Cat}_{\infty}^{\mathrm{perf}})$ the $\infty$-category of $2$-rings and symmetric monoidal exact functors. Recall that a $2$-ring is an essentially small, idempotent complete, symmetric monoidal and stable $\infty$-category whose tensor product is exact in each variable.
  7. (g)
    We denote by $\operatorname{CAlg}(\mathrm{Pr}^{\mathrm{L}}_{\mathrm{st}})$ the $\infty$-category of presentably symmetric monoidal stable $\infty$-categories and symmetric monoidal left adjoints.
  8. (h)
    We denote by $\mathrm{loc}_{\otimes}$ (resp. $\operatorname{thick}_{\otimes}$) the localizing (resp. thick) tensor ideal generated by a given class of objects.
  9. (i)
    For a finite group $G$, we let $\operatorname{Sp}_{G}$ denote the $\infty$-category of genuine $G$-spectra.
  10. (j)
    Given $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$, we write $\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)$ for the $\infty$-category of $R$-modules in $\operatorname{Sp}_{G}$ and $\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega}$ for its subcategory of compact objects. In the special case $G=e$, we know that the compact objects are equal to the perfect $R$-modules and so we will write $\operatorname{Perf}(R)$ instead of $\operatorname{Mod}_{\operatorname{Sp}}(R)^{\omega}$.
2. Preliminaries
In this section we fix some notation and recall a few key results that we will use throughout the paper.
Consider a symmetric monoidal stable $\infty$-category $\mathscr{C}$ and let $A\in\operatorname{CAlg}(\mathscr{C})$ be a commutative algebra object. Recall from [Bal11] that $A$ is separable if the multiplication map $A\otimes A\to A$ admits an $A$-bimodule section. We will write $\operatorname{CAlg}^{\operatorname{sep}}(\mathscr{C})$ for the full subcategory of commutative algebra objects $\operatorname{CAlg}(\mathscr{C})$ spanned by the separable algebras. It is immediate from the definition that any separable commutative algebra in $\mathscr{C}$ induces a separable commutative algebra object in the homotopy category $\operatorname{Ho}(\mathscr{C})$. The following result gives us a converse.
Theorem 2.1 (cf. [Ram23, Theorem B]).
Let $\mathscr{C}$ be a symmetric monoidal stable $\infty$-category, and let $A$ be an algebra in $\operatorname{Ho}(\mathscr{C})$ which is separable and commutative. Then $A$ admits an essentially unique lift $\widetilde{A}\in\operatorname{CAlg}(\mathscr{C})$ (and $\tilde{A}$ is separable). Furthermore, for any $R\in\operatorname{CAlg}(\mathscr{C})$, the canonical map
$\operatorname{Map}_{\operatorname{CAlg}(\mathscr{C})}(\widetilde{A},R)\xrightarrow{\sim}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Ho}(\mathscr{C}))}(A,R)$
is an equivalence.
A remarkable consequence of the previous result is that for any symmetric monoidal stable $\infty$-category $\mathscr{C}$, the $\infty$-category $\operatorname{CAlg}^{\operatorname{sep}}(\mathscr{C})$ is in fact a $1$-category.
Notation 2.2.
Let $\mathcal{C}\mathcal{S}\mathrm{ep}(-)$ denote the composite functor $\operatorname{CAlg}^{\operatorname{sep}}(-)^{\simeq}$ which takes the maximal groupoid of the subcategory of separable commutative algebra.
We recall the following result from  [Ram23, Corollary 3.24], which in a slightly different form also appears in [NP24, Proposition 6.3].
Proposition 2.3.
The functor
$\operatorname{CAlg}^{\operatorname{sep}}(-)\colon\operatorname{CAlg}(\operatorname{Cat}_{\infty}^{\mathrm{st}})\to\operatorname{Cat}_{\infty}$
preserves limits. Consequently, $\mathcal{C}\mathcal{S}\mathrm{ep}(-)$ also preserves limits.
3. Homotopy $G$-fixed points and separable algebras
In this section we will work with $\infty$-categories which are equipped with an action of a finite group. We start by recalling the definition of the homotopy $G$-fixed points of an $\infty$-category with $G$-action, and list various equivariant preliminaries regarding these categories. We then explain how, under the additional assumption that the $\infty$-category is suitably symmetric monoidal, the cotensor functor lifts to a functor taking values in a homotopy $G$-fixed points category.
Definition 3.1.
Give a group $G$ and $\mathscr{C}\in\operatorname{Fun}(BG,\operatorname{Cat}_{\infty})$, we can form
$\mathscr{C}^{hG}:=\lim_{BG}\mathscr{C}\in\operatorname{Cat}_{\infty}.$
Example 3.2.
If $G$ acts trivially on $\mathscr{C}$, then $\mathscr{C}^{hG}\simeq\operatorname{Fun}(BG,\mathscr{C})$.
Remark 3.3.
If $\mathscr{C}\in\operatorname{Fun}(BG,\Pr^{L}_{\mathrm{s}t})$, then the endomorphism ring of the unit object in $\mathscr{C}^{hG}$ is given by
$\operatorname{CAlg}\ni\operatorname{Map}_{\mathscr{C}^{hG}}(\mathbb{1},\mathbb{1})\simeq\lim_{BG}\operatorname{Map}_{\mathscr{C}}(\mathbb{1},\mathbb{1})\simeq\operatorname{Map}_{\mathscr{C}}(\mathbb{1},\mathbb{1})^{hG}.$
We record the following result from [NPR24, Proposition 6.7] which in the case $G$ acts trivially on $\mathscr{C}$, states the familiar fact that the restriction functor admits a biadjoint (induction and coinduction).
Lemma 3.4.
Let $G$ be a finite group and $\mathscr{C}\in\operatorname{Fun}(BG,\operatorname{Cat}_{\infty})$ whose underlying $\infty$-category is additive. For any subgroup $H\subseteq G$, the symmetric monoidal restriction functor
$\operatorname{res}_{H}^{G}\colon\mathscr{C}^{hG}\to\mathscr{C}^{hH}$
admits a left and a right adjoint, denoted respectively by $G\circledast_{H}(-)$ and $(-)^{\circledast_{H}G}$, and these two functors are equivalent. Furthermore, these adjunctions preserve dualizable objects.
Remark 3.5.
We observe that we always have canonical projection maps
$(s)^{\circledast_{H}G}\otimes t\to(s\otimes\operatorname{res}^{G}_{H}t)^{\circledast_{H}G}$
for all $t$ and $s$. Moreover, it is formal that these projection maps are equivalences if $t$ is dualizable, see for instance [NPR24, Lemma 3.8].
Remark 3.6.
We have a Mackey formula, namely for $H,K\subseteq G$, there is an equivalence
$\operatorname{res}_{H}^{G}(G\circledast_{K}-)\simeq\bigoplus_{g\in H\backslash G/K}H\circledast_{H\cap K^{g}}\operatorname{res}_{H\cap K^{g}}^{K}(-)$
where $K^{g}=gKg^{-1}$, and the restriction from $K$ to $H\cap K^{g}$ is along the conjugation map $H\cap K^{g}\subset K^{g}\cong K$. This is a consequence of [CLL23, Lemma 4.9] with $S=T=BG$ and $\mathcal{C}$ the $BG$-category given by
[…]
The reference shows that this is $BG$-cocomplete so in particular, the Beck-Chevalley transformation associated to the pullback in $\mathscr{S}_{/BG}\simeq\mathrm{PSh}(BG)$
${\coprod_{g\in H\backslash G/K}G/(H\cap K^{g})}$${BH}$${BK}$${BG}$
is an equivalence. One also checks that the limit-preserving extension of ()(*)( ∗ ) to $\mathscr{S}_{/BG}$ sends $BH\to BG$ to $\operatorname{Mod}(A)^{hH}$.
For example, when $K=e$ is the trivial subgroup, we find
$\operatorname{res}_{H}^{G}(G\circledast_{e}-)\simeq\bigoplus_{H\backslash G}(H\circledast_{e}-).$
In the following construction we explain how cotensoring defines a symmetric monoidal functor of the form $\operatorname{Fin}^{\operatorname{op}}\to\operatorname{CAlg}(\mathscr{C})$, $X\mapsto\mathbb{1}^{X}:=\prod_{X}\mathbb{1}$. This construction will be natural in symmetric monoidal $\infty$-categories $\mathscr{C}$ with finite products in which the tensor product preserves finite products in each variable, and finite product preserving symmetric monoidal functors between them.
Construction 3.7.
Let $\operatorname{Cat}_{\infty}^{\Pi}$ be the $\infty$-category of $\infty$-categories with finite products and finite product-preserving functors between them. It acquires a Lurie tensor product $\otimes$, following a dual construction to [Lur17, Section 4.8.1]. That is, $\mathscr{D}\otimes\mathscr{E}$ is such that finite product-preserving functors out of it into $\mathscr{F}$ are the same as functors $\mathscr{D}\times\mathscr{E}\to\mathscr{F}$ that preserve finite products in each variable separately.
As in [Lur17, Remark 4.8.1.9], commutative algebras in $(\operatorname{Cat}^{\Pi}_{\infty},\otimes)$ are exactly symmetric monoidal $\infty$-categories with finite products where the tensor product commutes with finite products in each variable. By design, the internal hom in $(\operatorname{Cat}^{\Pi}_{\infty},\otimes)$ is given by the category of finite product-preserving functors, and so $\operatorname{Fin}^{\operatorname{op}}$ is the unit for this symmetric monoidal structure since it is freely generated by the point under finite products [^1] Note that the finite products in $\operatorname{Fin}^{\operatorname{op}}$ are finite coproducts in $\operatorname{Fin}$..
In particular, $\operatorname{Fin}^{\operatorname{op}}$ is canonically and uniquely a commutative algebra in this category, and since the cocartesian symmetric monoidal structure [^2] Given by cartesian products in $\operatorname{Fin}$. is such a commutative algebra structure, it must be the relevant one. Thus, there is a symmetric monoidal, product-preserving functor $\operatorname{Fin}^{\operatorname{op}}\to\mathscr{C}$, natural in $\mathscr{C}\in\operatorname{CAlg}((\operatorname{Cat}^{\Pi}_{\infty},\otimes))$, where $\operatorname{Fin}^{\operatorname{op}}$ is equipped with the cocartesian symmetric monoidal structure. Since the forgetful functor induces a symmetric monoidal equivalence $\operatorname{CAlg}(\operatorname{Fin}^{\operatorname{op}})\simeq\operatorname{Fin}^{\operatorname{op}}$, this further induces a natural symmetric monoidal functor
$\mathbb{1}^{\bullet}\colon\operatorname{Fin}^{\operatorname{op}}\to\operatorname{CAlg}(\mathscr{C}).$
Finally, we note that for any finite set $X$, the commutative algebra $\mathbb{1}^{X}$ is a finite product of the unit object, so separable. Thus, the above functor lands naturally in $\operatorname{CAlg}^{\operatorname{sep}}(\mathscr{C})$. If $\mathscr{C}$ is semi-additive, which it will be in the rest of the paper, it is also dualizable, and thus the functor lands in $\operatorname{CAlg}^{\operatorname{sep}}(\mathscr{C}^{\operatorname{dual}}).$
Informally, pullback along the diagonal of a finite set $X$ is a separable, commutative algebra structure
$\mathbb{1}_{\mathscr{C}}^{X}\otimes\mathbb{1}_{\mathscr{C}}^{X}\simeq\mathbb{1}_{\mathscr{C}}^{X\times X}\to\mathbb{1}_{\mathscr{C}}^{X},$
natural in $\mathscr{C}\in\operatorname{CAlg}((\operatorname{Cat}_{\infty}^{\Pi},\otimes))$, and satisfying $\mathbb{1}_{\mathscr{C}}^{X\coprod Y}\simeq\mathbb{1}_{\mathscr{C}}^{X}\times\mathbb{1}_{\mathscr{C}}^{Y}$.
Construction 3.8.
Consider $\mathscr{C}\in\operatorname{Fun}(BG,\operatorname{CAlg}(\operatorname{Cat}_{\infty}^{\Pi}))$ which is in addition semi-additive. The naturality of 3.7 endows the functor
$\mathbb{1}^{\bullet}\colon\operatorname{Fin}^{\operatorname{op}}\to\operatorname{CAlg}^{\operatorname{sep}}(\mathscr{C}^{\operatorname{dual}})$
with a $G$-equivariant structure with respect to the trivial $G$-action on the source, and the canonical action coming from $\mathscr{C}$ on the target. By taking $G$-homotopy fixed points we obtain the symmetric monoidal and finite product-preserving functor
$(\mathbb{1}^{\bullet})^{hG}\colon\operatorname{Fin}_{G}^{\operatorname{op}}\to\operatorname{CAlg}^{\operatorname{sep}}(\mathscr{C}^{\operatorname{dual}})^{hG}.$
By construction, for any subgroup $H\subseteq G$ there is a natural comparison square:
[…]
which is vertically right adjointable by [NPR24, Lemma 6.9].
4. Stable module categories and the Tate construction
In this section we work with a commutative algebra $A\in\operatorname{CAlg}$ equipped with an action of a finite group $G$ and define three main categories
$\operatorname{Perf}(A)^{hG}\quad\mathrm{StMod}^{\mathrm{fin}}(AG)\quad\operatorname{Perf}(A)^{tG}$
which we will use throughout the paper. We start off with the first category.
Definition 4.1.
Let $G$ be a group and consider $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. The category of $A$-modules inherits a $G$-action, yielding a functor
$\operatorname{Mod}(A)\colon BG\to\mathrm{Pr}^{\mathrm{L}}_{\mathrm{st}}.$
We can then form $\operatorname{Mod}(A)^{hG}$ which we think of as the category of $A$-modules endowed with a semi-linear $G$-action. We note that the functor $\operatorname{Mod}(A)$ naturally lifts to $\operatorname{CAlg}(\mathrm{Pr}^{\mathrm{L}}_{\mathrm{st}})$ and so canonically $\operatorname{Mod}(A)^{hG}\in\operatorname{CAlg}(\mathrm{Pr}^{\mathrm{L}}_{\mathrm{st}})$. To connect this with the set up of the previous section, recall that the forgetful functor $\mathrm{Pr}^{\mathrm{L}}_{\mathrm{st}}\to\operatorname{Cat}_{\infty}$ preserves and creates limits, see proof of [NP24, Lemma 2.6].
Remark 4.2.
Using that taking dualizable objects commutes with limits (see [Lur17, Proposition 4.6.1.11]), we see that
$\left(\operatorname{Mod}(A)^{hG}\right)^{\mathrm{dual}}\simeq\left(\operatorname{Mod}(A)^{\mathrm{dual}}\right)^{hG}\simeq\operatorname{Perf}(A)^{hG}.$
Notation 4.3.
Let $G$ be a finite group and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. By applying 3.8 to $\mathscr{C}=\operatorname{Mod}(A)$ and using Proposition 2.3, we see that the cotensor functor extends to a functor
$A^{\bullet}\colon\operatorname{Fin}_{G}^{\operatorname{op}}\to\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(A)^{hG}),\quad X\mapsto A^{X}.$
By the discussion in the cited construction, for any subgroup $H\subseteq G$, there is an equivalence of commutative algebras
$A^{G/H}\simeq(\operatorname{res}^{G}_{H}A)^{\circledast_{H}G}.$
These commutative algebras $A^{X}$ migrate to any symmetric monoidal stable $\infty$-category receiving $\operatorname{Perf}(A)^{hG}$, like the Tate category of 4.12.
We next introduce the stable module category following [Kra20].
Notation 4.4.
Let $G$ be a finite group and $A\in\operatorname{CAlg}$. We set
$\mathrm{StMod}^{\mathrm{fin}}(AG):=\frac{\operatorname{Fun}(BG,\operatorname{Perf}(A))}{\operatorname{thick}_{\otimes}\langle A^{G/e}\rangle}\in\operatorname{Cat}_{\infty}^{\mathrm{st}}$
and refer to it as the stable module category of $G$ with coefficients in $A$. We also write
$\pi\colon\operatorname{Fun}(BG,\operatorname{Perf}(A))\to\mathrm{StMod}^{\mathrm{fin}}(AG)$
for the quotient functor.
Example 4.5.
If $A=k$ is a field, then $\operatorname{Ho}(\mathrm{StMod}^{\mathrm{fin}}(kG))$ is equivalent to the stable module category of $kG$ which appears in modular representation theory. This follows from the fact that Verdier quotients commute with taking homotopy categories and [Ric89, Theorem 2.1].
Remark 4.6.
We warn the reader that the stable module category is not idempotent complete in general. Our definition agrees with that of Mathew [Mat15] up to idempotent completion. See [Kra20, Remark 4.3] for more details.
Remark 4.7.
As the stable module category is defined as a Verdier quotient by a thick ideal, it inherits a symmetric monoidal structure from $\operatorname{Fun}(BG,\operatorname{Perf}(A))$, see for example [NS18, Theorem I.3.6]. It follows that $\mathrm{StMod}^{\mathrm{fin}}(AG)\in\operatorname{CAlg}(\operatorname{Cat}_{\infty}^{\mathrm{st}})$ in such a way that the canonical functor $\pi\colon\operatorname{Fun}(BG,\operatorname{Perf}(A))\to\mathrm{StMod}^{\mathrm{fin}}(AG)$ is symmetric monoidal.
Remark 4.8.
Consider $X,Y\in\operatorname{Fun}(BG,\operatorname{Perf}(A))$. Then by [Kra20, Lemma 4.2] we find that
$\operatorname{Map}_{\mathrm{StMod}^{\mathrm{fin}}(AG)}(X,Y)\simeq\operatorname{Map}_{\operatorname{Perf}(A)}(X,Y)^{tG}\simeq\left(\mathbb{D}X\otimes Y\right)^{tG}.$
As in 3.3, this identifies the endomorphism ring of the unit of $\mathrm{StMod}^{\mathrm{fin}}(AG)$ with $A^{tG}\in\operatorname{CAlg}$.
We record the following result for future reference.
Lemma 4.9.
Let $G$ be a finite group of order $g$, and suppose that $A\in\operatorname{CAlg}$ is bounded below. Then:
  • (a)
    The canonical map $A^{\wedge}_{g}\to\prod_{0<p|g}A_{p}^{\wedge}$, from the $g$-completion to the product of the $p$-completions, is an equivalence.
  • (b)
    The functor induced by base-change
    $\beta\colon\mathrm{StMod}^{\mathrm{fin}}(AG)\to\mathrm{StMod}^{\mathrm{fin}}(A^{\wedge}_{g}G)\simeq\prod_{p\mid g}\mathrm{StMod}^{\mathrm{fin}}(A^{\wedge}_{p}G)$
    is fully faithful.
Proof.
Part (a) is true more generally for an arbitrary spectrum $X$ in place of $A$. For any abelian group $C$ and coprime integers $a,b\geq 1$ both of the group homomorphisms
$C/ab\stackrel{{\scriptstyle\mathrm{red}}}{{\longrightarrow}}C/a\oplus C/b\,\,\mbox{ and }\,\,C[ab]\stackrel{{\scriptstyle(\cdot b,\cdot a)}}{{\longrightarrow}}C[a]\oplus C[b]$
are isomorphisms. The map of fiber sequences
${X}$${X}$${X/ab}$${X}$${X}$${X/a}$$\scriptstyle{b}$$\scriptstyle{ab}$$\scriptstyle{\mathrm{id}}$$\scriptstyle{a}$
induces a map of short exact sequences
00{0}${(\pi_{*}(X))/ab}$${\pi_{*}(X/ab)}$${(\pi_{*}(\Sigma X))[ab]}$00{0}00{0}${(\pi_{*}(X))/a}$${\pi_{*}(X/a)}$${(\pi_{*}(\Sigma X))[a]}$${0.}$$\scriptstyle{\mathrm{red}}$$\scriptstyle{\cdot b}$
Reversing the roles of $a$ and $b$ induces a similar map of short exact sequences which combine into
00{0}${(\pi_{*}(X))/ab}$${\pi_{*}(X/ab)}$${(\pi_{*}(\Sigma X))[ab]}$00{0}00{0}${(\pi_{*}(X))/a\oplus(\pi_{*}(X))/b}$${\pi_{*}(X/a)\oplus\pi_{*}(X/b)}$${(\pi_{*}(\Sigma X))[a]\oplus(\pi_{*}(\Sigma X))[b]}$${0.}$$\scriptstyle{\mathrm{red}}$$\scriptstyle{(\cdot b,\cdot a)}$
Since the outer maps are isomorphisms by our initial observation, the map $X/ab\to X/a\times X/b$ is also an equivalence. One then obtains the claim by induction on the number of prime factors of $g$ which obviously starts with the case that $g$ is a prime power. In the induction step, one has $g=p^{n}g^{\prime}$ for some prime $p$ not dividing $g^{\prime}$, and can apply the above with $a=p^{nk}$ and $b=(g^{\prime})^{k}$ for an arbitrary $k\geq 1$, obtaining
$X/p^{nk}\times X/(g^{\prime})^{k}\simeq X/g^{k}.$
Passage to the limit over $k$ yields $X_{p}^{\wedge}\times X_{g^{\prime}}^{\wedge}\simeq X_{g}^{\wedge}$, as desired.
The proof of (b) is essentially [Kra20, Lemma 5.1]. We recall the argument here for completeness. We only show that the base-change map is fully faithful, the fact that the stable module category splits as a product is an easy consequence of the fact that $A_{g}^{\wedge}$ itself splits. For this, by 4.8, we need to check that the map
(4.10) $(\mathbb{D}X\otimes_{A}Y)^{tG}\to((A_{g}^{\wedge}\otimes_{A}\mathbb{D}X)\otimes_{A_{g}^{\wedge}}(A_{g}^{\wedge}\otimes_{A}Y))^{tG}\simeq(A_{g}^{\wedge}\otimes_{A}\mathbb{D}X\otimes_{A}Y)^{tG}$
is an equivalence for all $X,Y\in\operatorname{Fun}(BG,\operatorname{Perf}(A))$. Recall that by [NS18, Lemma I.2.9], the Tate construction of any bounded below spectrum is $g$-complete  [^3] The cited result is stated for $G=C_{p}$ but the same proof applies to any finite group $G$.. Thus if $A$ is bounded below, this holds for any perfect $A$-module such as $\mathbb{D}X\otimes_{A}Y$.
In particular, the modules appearing in (4.10) are $g$-complete, and so to verify that the map is an equivalence, we can do so after tensoring with $A^{tG}/g$. Using exactness of the Tate construction, this reduces to checking that the reduction mod $g$ of $A$ and $A^{\wedge}_{g}$ agree, which is clear. ∎
Remark 4.11.
An alternative argument for the key point of Lemma 4.9(a) is to recall that every stable category is tensored over finite spectra, so one needs to see that for coprime integers $a$ and $b$, the map
$\mathbb{S}/ab\longrightarrow\mathbb{S}/a\times\mathbb{S}/b$
constructed there is an equivalence. Since these spectra are connective, it suffices by the Hurewicz theorem to see that the above map in an equivalence after tensoring with $\mathbb{Z}$, which is true by the Chinese remainder theorem.
Finally we introduce the last category of interest and explain how it relates to the stable module category.
Definition 4.12.
Let $G$ be a finite group and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$.
  • (a)
    We define the homotopy $G$-orbits of $\operatorname{Perf}(A)$ to be
    $\operatorname{Perf}(A)_{hG}:=\operatorname{colim}_{BG}\operatorname{Perf}(A)$
    where the colimit is calculated in $\operatorname{Cat}_{\infty}^{\mathrm{perf}}$.
  • (b)
    There is a canonical comparison map between the colimit and the limit
    $\mathrm{Nm}_{G}\colon\operatorname{Perf}(A)_{hG}\to\operatorname{Perf}(A)^{hG}$
    which identifies with the norm map. We define the $G$-Tate of $\operatorname{Perf}(A)$ to be
    $\operatorname{Perf}(A)^{tG}:=\operatorname{cof}(\operatorname{Perf}(A)_{hG}\xrightarrow{\mathrm{Nm}_{G}}\operatorname{Perf}(A)^{hG})\in\operatorname{Cat}_{\infty}^{\mathrm{perf}};$
    the cofibre of the norm map in $\operatorname{Cat}_{\infty}^{\mathrm{perf}}$. Note that there is a canonical functor $q\colon\operatorname{Perf}(A)^{hG}\to\operatorname{Perf}(A)^{tG}$.
We now record a more explicit description of the $G$-Tate of $\operatorname{Perf}(A)$. We denote by $(-)^{\natural}$ the idempotent completion functor.
Lemma 4.13.
Let $G$ be a finite group and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. Then
$\operatorname{Perf}(A)^{tG}\simeq\left(\frac{\operatorname{Perf}(A)^{hG}}{\operatorname{thick}_{\otimes}\langle A^{G/e}\rangle}\right)^{\natural}.$
Proof.
Since cofibres in $\operatorname{Cat}_{\infty}^{\mathrm{perf}}$ are given by the idempotent completion of the Verdier quotient, we need to check that the image of $\mathrm{Nm}_{G}$ is $\operatorname{thick}_{\otimes}\langle A^{G/e}\rangle$. By [CMNN24, Proposition 2.18] taking $\mathcal{A}=\operatorname{Perf}(A)\in\operatorname{Fun}(BG,\operatorname{Cat}_{\infty}^{\mathrm{perf}})$ there, we know that the image of the norm map agrees with $\operatorname{thick}(A^{G/e})$ which is automatically an ideal by an application of the projection formula (see 3.5). ∎
Example 4.14.
Suppose that $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ with trivial $G$-action. Then $\operatorname{Perf}(A)^{tG}$ is the idempotent completion of $\mathrm{StMod}^{\mathrm{fin}}(AG)$.
Remark 4.15.
We note that $\operatorname{Perf}(A)^{tG}$ acquires a symmetric monoidal structure from $\operatorname{Perf}(A)^{hG}$ making the functor $q\colon\operatorname{Perf}(A)^{hG}\to\operatorname{Perf}(A)^{tG}$ into a symmetric monoidal functor, see discussion after [Mat16, Definition 2.16]. It follows that $\operatorname{Perf}(A)^{tG}\in\operatorname{CAlg}(\operatorname{Cat}_{\infty}^{\mathrm{perf}})$ and so $\operatorname{Ind}(\operatorname{Perf}(A)^{tG})\in\operatorname{CAlg}(\mathrm{Pr}^{\mathrm{L}}_{\mathrm{st}})$.
Lemma 4.16.
Let $H$ be a subgroup of a finite group $G$ and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. Then the restriction-coinduction adjunction descends to the Tate categories
$\operatorname{res}^{G}_{H}\colon\operatorname{Perf}(A)^{tG}\rightleftarrows\operatorname{Perf}(A)^{tH}:(-)^{\circledast_{H}G}$
Furthermore, these last functors induce adjoint pairs at the level of commutative algebra objects.
Proof.
We claim that $\operatorname{res}_{H}^{G}A^{G/e}\in\operatorname{thick}_{\otimes}\langle A^{H/e}\rangle$. From this and the universal property of the Verdier quotient we see that the symmetric monoidal functor $\operatorname{res}^{G}_{H}\colon\operatorname{Perf}(A)^{hG}\to\operatorname{Perf}(A)^{hH}$ descends to a symmetric monoidal functor between the Tate categories. Since $A^{G/e}\simeq(A^{H/e})^{\circledast_{H}G}$ also the right adjoint passes to the Tate categories. It also follows that $(-)^{\circledast_{H}G}$ acquires a lax symmetric monoidal structure so the adjunction descends to commutative algebra objects. Therefore it is only left to prove the claim. By Lemma 3.4 we can instead verify that $\operatorname{res}_{H}^{G}(G\circledast_{e}A)\in\operatorname{thick}_{\otimes}\langle H\circledast_{e}A\rangle$. By the Mackey formula (cf. 3.6), $\operatorname{res}_{H}^{G}(G\circledast_{e}A)\simeq\bigoplus_{G/H}(H\circledast_{e}A)$ and is therefore indeed in $\operatorname{thick}_{\otimes}\langle H\circledast_{e}A\rangle$. ∎
5. The indecomposable condition
Let $\operatorname{\mathscr{C}}$ be a symmetric monoidal $\infty$-category with unit object $\mathbb{1}$. Recall that an idempotent of a commutative algebra $A\in\operatorname{CAlg}(\operatorname{\mathscr{C}})$ is by definition an idempotent element of the discrete ring $\pi_{0}(A):=\pi_{0}(\operatorname{Map}_{\operatorname{\mathscr{C}}}(\mathbb{1},A))$. We then say that $A$ is indecomposable if $A$ is nonzero and if it does not decompose as a product of two nonzero rings. This is equivalent to $\pi_{0}(A)$ being an indecomposable ring. See [NP24, Section 3] for more details.
Remark 5.1.
Let $\operatorname{\mathscr{C}}$ be a symmetric monoidal $\infty$-category and consider $R,S\in\operatorname{CAlg}(\operatorname{\mathscr{C}})$ with $S$ indecomposable. For any decomposition $R=\prod_{i=1}^{n}R_{i}$ into a finite product, we obtain
$\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(R,S)=\coprod_{i=1}^{n}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(R_{i},S).$
We next introduce one of our conditions which we require to prove our main result.
Definition 5.2.
Let $G$ be a finite group and consider $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. We say that $A$ satisfies the indecomposable condition (in short IC) if the commutative algebras $A^{hG}$ and $A^{tG}$ are indecomposable.
We next give a few examples of commutative algebras satisfying this condition.
Proposition 5.3.
Let $G$ be a finite group and let $A\in\operatorname{Fun}(BG,\operatorname{CAlg}^{\heartsuit})$ be a discrete commutative algebra with trivial $G$-action. Then $A$ satisfies IC if and only if $A$ and $A/|G|$ are indecomposable.
Proof.
Using the homotopy fixed points and Tate spectral sequences we see that
$\pi_{0}(A^{hG})=H^{0}(G,A)=A\quad\mathrm{and}\quad\pi_{0}(A^{tG})=\widehat{H}^{0}(G,A)=A/|G|.$
Example 5.4.
If $G$ is a nontrivial $p$-group, then the integers $\mathbb{Z}$ and any field $k$ of characteristic $p$ satisfy IC with respect to the trivial $G$-action.
Theorem 5.5.
Let $G$ be a finite group and let $\mathbb{S}_{G}\in\operatorname{Fun}(BG,\operatorname{CAlg})$ be the sphere spectrum with trivial $G$-action. Then $\mathbb{S}_{G}$ satisfies IC if and only if $G$ is a nontrivial $p$-group for some prime number $p$.
Proof.
We start by calculating $\pi_{0}(\mathbb{S}_{G}^{hG})$ and $\pi_{0}(\mathbb{S}_{G}^{tG})$. By the Segal conjecture, we have
$\pi_{0}(\mathbb{S}_{G}^{hG})=\pi_{0}(\mathbb{S}^{BG})=\widehat{A(G)}_{I}$
the completion of the Burnside ring of $G$ at the augmentation ideal $I$. For the Tate construction we look at the defining fiber sequence
[…]
Since $\pi_{-1}((\mathbb{S}_{G})_{hG})=0$ and $\pi_{0}((\mathbb{S}_{G})_{hG})=\mathbb{Z}$, we deduce that the ring homomorphism $\pi_{0}(\mathbb{S}_{G}^{hG})\to\pi_{0}(\mathbb{S}_{G}^{tG})$ is surjective with kernel generated by $N(1)=[G]\in\widehat{A(G)}_{I}$. Hence
$\pi_{0}(\mathbb{S}_{G}^{tG})=\widehat{A(G)}_{I}/([G]).$
Next we check when these two rings are indecomposable. To this end, recall that for any henselian pair $(R,I)$, the reduction $R\to R/I$ induced a bijection on idempotents, because the $R$-algebra $R[X]/(X^{2}-X)$ classifying idempotents is étale, see [Sta24, Tag 09XI]. This in particular applies to the $I$-adically complete ring $\widehat{A(G)}_{I}$. It also applies to the quotient ring $\widehat{A(G)}_{I}/([G]))$ as it is a finitely generated module over the complete Noetherian ring $\widehat{A(G)}_{I}$, hence also $I$-adically complete. We are thus reduced to checking when the two rings
$\widehat{A(G)}_{I}/I\quad\mathrm{and}\quad\widehat{A(G)}_{I}/(([G])),I)$
are indecomposable. Since the first quotient is $\mathbb{Z}$, and $[G]-|G|\in I$, we are reduced to checking when the two rings $\mathbb{Z}$ and $\mathbb{Z}/|G|$ are indecomposable. By the Chinese Remainder Theorem this holds if and only if $G$ is a nontrivial $p$-group for some prime $p$. ∎
Example 5.6.
For a finite group $G$, consider $KU\in\operatorname{Fun}(BG,\operatorname{CAlg})$, the spectrum of topological $K$-theory with trivial $G$-action. Then $\pi_{0}(KU^{hG})$ is the completion of the complex representation ring $RU(G)$ at the augmentation ideal $I$. The observation about henselian pairs in the previous proof shows that the quotient map
$RU(G)^{\wedge}_{I}\to RU(G)^{\wedge}_{I}/I\simeq\mathbb{Z}$
induces a bijection on idempotents. Thus $KU^{hG}$ is indecomposable for every finite group $G$. On the other hand [NS18, Example I.2.3.(iii)] shows that $KU^{tG}$ is not indecomposable if $G=C_{p^{2}}.$
6. The retraction condition
In this section we isolate the following condition:
Definition 6.1.
Let $G$ be a finite group and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$. We say that $A$ satisfies the retraction condition (in short RC) if for all $H\subseteq G$ such that $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{tG}$, we have $H=G$.
In applications, we will use the following consequence of RC.
Remark 6.2.
Suppose that $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfies RC. Then $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{hG}$ if and only if $H=G$. The “if” part is clear and for the “only if” part note that if $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{hG}$, then we can apply the canonical functor $\operatorname{Perf}(A)^{hG}\to\operatorname{Perf}(A)^{tG}$ to this retraction diagram to find, by RC, that $H=G$.
This condition lets us prove the following result.
Theorem 6.3.
Let $G$ be a finite group and let $X,Y$ be finite $G$-set. Suppose that the isotropy of $Y$ does not contain the trivial group. Let $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfy IC and RC. Then the canonical map
$\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Perf}(A)^{hG})}(A^{X},A^{Y})\xrightarrow{\sim}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{Perf}(A)^{tG})}(A^{X},A^{Y})$
is an equivalence.
Proof.
We note that if $G=e$, then $Y=\varnothing$ and so $A^{Y}=0$. The canonical map in the theorem is then an equivalence as source and target are contractible. So without loss of generality we can assume that $G$ is nontrivial. Firstly, let us treat the special case in which $Y=G/G$ so that $A^{Y}=A$. Decomposing $X$ into its orbits $X\simeq\coprod_{i}G/H_{i}$, we see that $A^{X}\simeq\prod_{i}A^{G/H_{i}}$. Note that the unit $A$ is indecomposable both in $\operatorname{Perf}(A)^{hG}$ and $\operatorname{Perf}(A)^{tG}$ by the assumption IC for $A$. Remark 5.1 then implies that the relevant mapping space of commutative algebras in either category is given by
$\operatorname{Map}(A^{X},A)\simeq\coprod_{i}\operatorname{Map}(A^{G/H_{i}},A).$
We are therefore reduced to showing that passage to the quotient category induces an equivalence on the mapping spaces $\operatorname{Map}(A^{G/H},A)$. Indeed, we claim these mapping spaces are simultaneously empty or contractible, depending on $H$: If $H\not=G$, then we claim that both mapping spaces are empty as there are not even any unital maps from $A^{G/H}$ to $A$. This follows from the fact that $A$ satisfies RC and 6.2. If $H=G$, then both mapping spaces are contractible as the unit $A=A^{G/G}$ is the initial commutative algebra in both categories. We have therefore proved the theorem in the case that $Y=G/G$.
For the general case, decompose $Y$ into orbits $Y\simeq\coprod_{i}G/H_{i}$, and note that by our assumption $H_{i}\not=e$ for all $i$. Then in both categories we find that
$\operatorname{Map}(A^{X},A^{Y})=\prod_{i}\operatorname{Map}(A^{X},A^{G/H_{i}})=\prod_{i}\operatorname{Map}(\operatorname{res}^{G}_{H_{i}}(A^{X}),A),$
where we used that $A^{G/H_{i}}$ is coinduced, see 4.3. Then the claim follows from the previous paragraph, observing that for each $i$, we have
$\operatorname{res}^{G}_{H_{i}}(A^{X})\simeq\prod_{j}A^{H/H_{ij}},$
where
$\operatorname{res}^{G}_{H_{i}}(X)\simeq\coprod_{j}H_{i}/H_{ij}$
is the decomposition of the $H_{i}$-space into orbits. (Note this argument uses that $A$ is indecomposable in the Tate category of each $H_{i}$). ∎
We now list some examples of commutative algebra objects which satisfies RC.
Lemma 6.4.
Let $G$ be a finite group and let $A\in\operatorname{CAlg}^{\heartsuit}$ be a $|G|$-torsion free discrete commutative ring in which no prime divisor of $|G|$ is invertible. Then $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ with trivial $G$-action satisfies RC.
Proof.
Suppose that $A$ is a retract of $A^{G/H}$ in $\operatorname{Perf}(A)^{tG}$ for some $H\subseteq G$. We need to show that $H=G$. Applying
$\pi_{0}\operatorname{Map}_{\operatorname{Perf}(A)^{tG}}(A,-)\colon\operatorname{Perf}(A)^{tG}\longrightarrow\operatorname{Mod}_{\pi_{0}(A^{tG})}(\operatorname{Ab})=\operatorname{Mod}_{A/|G|}(\operatorname{Ab})$
to a retraction diagram and using 3.3 and the fact that $A^{G/H}$ is a coinduced module, we find that the $A$-module $A/|G|$ is a retract of $A/|H|$. Note that $A/|H|$ is $|H|$-torsion, therefore so is $A/|G|$. It follows that $|H|\cdot 1_{A}=0$ mod $|G|$, so there is $a\in A$ such that $a[G:H]|H|=|H|\cdot 1_{A}$. The $|G|$- and hence $|H|$-torsion-freeness then implies that $a[G:H]=1$, so $[G:H]$ is invertible in $A$. By assumption, this forces $|H|=|G|$ and thus $H=G$. ∎
Lemma 6.5.
Let $G$ be a finite group and let $A\to B$ be a map in $\operatorname{Fun}(BG,\operatorname{CAlg})$. Then RC for $B$ implies RC for $A$.
Proof.
The naturality in $\mathscr{C}$ in 3.8 implies $B\otimes_{A}A^{G/H}\simeq B^{G/H}$ in $\operatorname{Perf}(B)^{tG}$. ∎
7. Separably closed commutative algebras
In this section we introduce the last condition required to prove our main result.
Definition 7.1.
We say that $A\in\operatorname{CAlg}$ is separably closed if the cotensor functor is an equivalence
$A^{(-)}\colon\operatorname{Fin}^{\operatorname{op}}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(A)).$
If $G$ is a finite group and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$, then we say that $A$ is separably closed if the underlying commutative algebra $A\in\operatorname{CAlg}$ is separably closed.
Let us give a few remarks on this definition.
Remark 7.2.
Any separably closed commutative algebra $A$ is necessarily indecomposable. Indeed using that retracts of separable algebras are separable, we see that a nontrivial decomposition of $A$ would give a nontrivial decomposition of $\ast\in\operatorname{Fin}$, and so a contradiction.
Remark 7.3.
We warn the reader that our definition of separably closed commutative algebra is more restrictive than the one given by Rognes in [Rog08, Definition 10.3.1]. This is because any Galois extension is in particular separable, see [Rog08, Lemma 9.1.2]. To see that our definition is strictly stronger, we fix a separably closed field $k$ of characteristic $p$, and consider $A=k^{tE}$ for an elementary abelian $p$-group $E$. Then $A$ is separably closed in the sense of Rognes by work of Mathew [Mat16, Theorem 9.9]. On the other hand by [Mat16, Theorem 2.30], the $\infty$-category $\operatorname{Perf}(k^{tE})$ identifies with the small stable module category of $kE$, and in there we know that there are separable algebras which are not just products of $k$, for instance $k[E/E_{0}]$ for $e\not=E_{0}\subset E$, as one checks by contemplating support in the Balmer spectrum.
Let us give a few examples.
Lemma 7.4.
Let $S\in\operatorname{CAlg}$ be separably closed and let $R:=\operatorname{infl}_{e}^{G}S\in\operatorname{CAlg}(\operatorname{Sp}_{G})$. Then, for all subgroups $K\subseteq G$, the geometric fixed points $\Phi^{K}R$ are separably closed.
Proof.
For every $K\subseteq G$, there is an equivalence $\Phi^{K}\operatorname{infl}_{e}^{G}S\simeq S$. ∎
Example 7.5.
For instance we can take $S=\mathbb{S}$ so that $R=\mathbb{S}_{G}$ or $S=\mathbb{Z}$ so that $R=\operatorname{infl}_{e}^{G}(\mathbb{Z})$. The fact that the rings $S$ are separably closed follows from Minkowski’s theorem, see [NP24, Proposition 10.5].
Next, we record a few consequences of this condition.
Lemma 7.6.
Let $G$ be a finite group and $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ separably closed. Then the functor from 4.3
$A^{(-)}\colon\operatorname{Fin}^{\operatorname{op}}_{G}\xrightarrow{\sim}\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(A)^{hG})$
is an equivalence.
Proof.
Our assumption that $A$ is separably closed ensures that
$A^{(-)}\colon\operatorname{Fin}^{\operatorname{op}}\xrightarrow{\sim}\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(A))$
is an equivalence. The claim then follows from this by taking $G$-homotopy fixed points. ∎
Notation 7.7.
Let $\mathscr{F}$ be a family of subgroups of a finite group $G$, i.e. a possibly empty set of subgroups of $G$, closed under sub-conjugation. We denote by $\operatorname{Fin}_{G}/\mathscr{F}$ the full subcategory of $\operatorname{Fin}_{G}$ spanned by those finite $G$-sets $X$ which admits a (possibly empty) decomposition into orbits $X\simeq\coprod_{i}G/H_{i}$ with $H_{i}\not\in\mathscr{F}$. We also denote by $\operatorname{Fin}_{G}^{\mathrm{free}}$ the category of finite free $G$-sets. To simplify notation we we will write
$\mathcal{F}\mathrm{in}_{G}$ $:=\operatorname{Fin}_{G}^{\simeq}$
$\mathcal{F}\mathrm{in}_{G}/\mathscr{F}$ $:=(\operatorname{Fin}_{G}/\mathscr{F})^{\simeq}$
$\mathcal{F}\mathrm{in}_{G}^{\mathrm{free}}$ $:=(\operatorname{Fin}_{G}^{\mathrm{free}})^{\simeq}$
using caligraphic letters to indicate the maximal groupoid underlying a category. We note that the above groupoids canonically carry a cocartesian monoidal structure turning them into symmetric monoidal groupoids, or equivalently, into $\mathbb{E}_{\infty}$-monoids in spaces.
Remark 7.8.
There is a symmetric monoidal (wrt disjoint union) equivalence of groupoids
(7.9) $\mathcal{F}\mathrm{in}_{G}/\mathscr{F}\simeq\prod_{(H)\not\in\mathscr{F}}\mathcal{F}\mathrm{in}_{W_{G}(H)}^{\mathrm{free}},\quad X\mapsto(\mathrm{Map}_{G}^{\mathrm{inj}}(G/H,X))_{(H)}$
where the product is taken over a set of representatives of $G$-conjugacy classes of subgroups $H\not\in\mathscr{F}$, and $\mathrm{Map}_{G}^{\mathrm{inj}}$ denotes the set of $G$-equivariant injective maps.
Recall the functor $q$ of passage to the Tate-category from 4.12.
Lemma 7.10.
Let $G$ be a finite group and assume $A\in\operatorname{Fun}(BG,\operatorname{CAlg})$ satisfies IC and RC. Then, there is a unique factorization
${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(A)^{hG})}$${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(A)^{tG})}$${\mathcal{F}\mathrm{in}^{\operatorname{op}}_{G}\simeq\prod_{(H)}(\mathcal{F}\mathrm{in}_{W_{G}(H)}^{\mathrm{free}})^{\operatorname{op}}}$${\prod_{e\not=(H)}(\mathcal{F}\mathrm{in}_{W_{G}(H)}^{\mathrm{free}})^{\operatorname{op}},}$$\scriptstyle{q}$$\scriptstyle{A^{(-)}}$$\scriptstyle{\mathrm{proj}}$$\scriptstyle{\mathrm{\iota}}$
and $\iota$ is fully faithful.
Proof.
Since the product in the bottom left corner is a product in the preadditive $\infty$-category of $\mathbb{E}_{\infty}$-monoids in spaces, it is also a coproduct (see [GGN15, Proposition 2.3]), and the indicated $\iota$ exists if and only if $q\circ A^{(-)}$ is zero on the factor corresponding to $(H)=e$. This follows from the fact that the top horizontal functor sends $A^{G/e}$ to a terminal object, see Lemma 4.13.
Uniqueness follows from the fact that the nullhomotopy is unique: the component of 00 in $\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(A)^{tG})$ is contractible because 00 is terminal. Alternatively, one can consider this as a square of semirings in spaces, where uniqueness is because projecting off a component amounts to inverting an idempotent.
Fully faithfulness of $\iota$ follows from Theorem 6.3 and the fact that fully faithfulness is preserved by the groupoid functor. ∎
8. Pullback decomposition of equivariant spectra
In this section we recall the pullback decomposition for equivariant spectra given in [NPR24, Section 7] which will be the key tool for classifying separable algebras in this setting. Here we will be brief and refer the interested reader to the above reference for more details. A similar pullback decomposition already appears in [Kra20, Theorem 3.11] for inflated commutative ring spectra; we refer the reader to the introduction of [NPR24] for a more in depth discussion of the relation between these two results.
Fix a finite group $G$ and a family $\mathscr{F}$ of subgroups of $G$. The case $\mathscr{F}=\varnothing$ is allowed here. Given a family $\mathscr{F}$ and a subset $S$ of the set of subgroups of $G$, we let $\mathscr{F}\cup S$ denote the smallest family containing $\mathscr{F}$ and $S$. We will exclusively be concerned with the case $S=\{K\}$ where $K\subseteq G$ is a subgroup such that $K\notin\mathscr{F}$, but for all proper subgroups $K^{\prime}\subsetneq K$ we have $K^{\prime}\in\mathscr{F}$.
We denote by $\operatorname{Sp}_{G}/\mathscr{F}$ the $\infty$-category of genuine $G$-spectra with isotropy outside $\mathscr{F}$. Specifically, this is the smashing localization of $\operatorname{Sp}_{G}$ corresponding to the idempotent commutative algebra $\widetilde{E}\mathscr{F}$ where we have omitted the suspension spectrum functor from the notation. The canonical map $\widetilde{E}\mathscr{F}\to\widetilde{E}(\mathscr{F}\cup K)$ induces a further localization
$\operatorname{Sp}_{G}/\mathscr{F}\to\operatorname{Sp}_{G}/\mathscr{F}\cup{K}$
which annihilates $\widetilde{E}\mathscr{F}\otimes G/K_{+}$.
Notation 8.1.
Fix $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$ and write
$\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)\quad\mathrm{and}\quad\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}$
for the $\infty$-category of $R$-modules in $\operatorname{Sp}_{G}$ and $\operatorname{Sp}_{G}/\mathscr{F}$ respectively. More concretely, we have $\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\simeq\operatorname{Mod}_{\operatorname{Sp}_{G}}(R\otimes\widetilde{E}\mathscr{F})$. We denote by
$L_{\mathscr{F}\cup K}\colon\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\to\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K$
the induced smashing localization which annihilates $R\otimes\widetilde{E}\mathscr{F}\otimes G/K_{+}$.
Let $\Phi^{K}\colon\operatorname{Sp}_{G}\to\operatorname{Sp}$ denote the $K$-geometric fixed points functor, and let
$\Phi^{K}_{R}\colon\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)\to\operatorname{Mod}(\Phi^{K}(R))$
denote its $R$-linear extension. Recall that $\Phi^{K}(R)$ is acted on by the Weyl-group $W_{G}(K)$ of $K$ in $G$, allowing us to form
$\operatorname{Perf}(\Phi^{K}(R))^{hW_{G}(K)}.$
As discussed in [NPR24, Proposition 7.7], the $R$-linear geometric $K$-fixed points functor factors through some refined variant $\widetilde{\Phi}_{R}^{K}$ which remembers the Weyl-group action, as follows:
$\Phi_{R}^{K}\colon\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)\xrightarrow{\widetilde{\Phi}_{R}^{K}}\operatorname{Mod}(\Phi^{K}(R))^{hW_{G}(K)})\xrightarrow{\mathrm{fgt}}\operatorname{Mod}(\Phi^{K}(R)).$
Moreover, using that $\widetilde{\Phi}_{R}^{K}$ annihilates the kernel of $L_{\mathscr{F}\cup K}$, we get an induced functor
$(\overline{\Phi}^{K}_{R})^{\omega}\colon(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K)^{\omega}\to\operatorname{Perf}(\Phi^{K}R)^{tW_{G}(K)}.$
Finally, recall the functor $q$ from 4.12.
Theorem 8.2 ([NPR24, Corollary 7.12]).
Let $G$ be a finite group and $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$. Let $\mathscr{F}$ be a family of subgroups of $G$ and fix a subgroup $K\subseteq G$ such that $K\notin\mathscr{F}$, but for all proper subgroups $K^{\prime}\subsetneq K$ we have $K^{\prime}\in\mathscr{F}$. Then the following is a pullback of symmetric monoidal stable $\infty$-categories:
[…]
Recall the functors $\pi$ and $\beta$ from 4.4 and Lemma 4.9.
Corollary 8.3.
Let $G$ be a finite group, $\mathscr{F}$ a family of subgroups of $G$ and let $K\subseteq G$ be a subgroup such that $K\not\in\mathscr{F}$ and $K^{\prime}\in\mathscr{F}$ for all $K^{\prime}\subsetneq K$. Suppose we are given $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$ such that $\Phi^{K}(R)$ is bounded below and $W_{G}(K)$ acts trivially on $\Phi^{K}(R)$. Then the following is a pullback of symmetric monoidal stable $\infty$-categories
${(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}}$${(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K)^{\omega}}$${\operatorname{Fun}(BW_{G}(K),\operatorname{Perf}(\Phi^{K}(R)))}$${\prod_{0<p\mid|W_{G}(K)|}\mathrm{StMod}^{\mathrm{fin}}((\Phi^{K}R)_{p}^{\wedge}W_{G}(K)).}$$\scriptstyle{L_{\mathscr{F}\cup K}^{\omega}}$$\scriptstyle{(\widetilde{\Phi}_{R}^{K})^{\omega}}$$\scriptstyle{\beta\circ(\overline{\Phi}^{K}_{R})^{\omega}}$$\scriptstyle{\beta\circ\pi}$
Proof.
By [NPR24, Corollary 7.13], the following is a pullback square
${(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}}$${(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K)^{\omega}}$${\operatorname{Fun}(BW_{G}(K),\operatorname{Perf}(\Phi^{K}(R)))}$${\mathrm{StMod}^{\mathrm{fin}}((\Phi^{K}R)W_{G}(K)).}$$\scriptstyle{L_{\mathscr{F}\cup K}^{\omega}}$$\scriptstyle{(\widetilde{\Phi}_{R}^{K})^{\omega}}$$\scriptstyle{(\overline{\Phi}_{R}^{K})^{\omega}}$$\scriptstyle{\pi}$
We observe that the above square differs from the one in the theorem by the bottom right corner. However, by Lemma 4.9, the functor $\beta$ relating the two vertices is fully faithful, and so the pullbacks of the two squares are equivalent. ∎
9. The classification result
Given a finite $G$-set $X$, we can form the commutative algebra $\mathbb{D}(\Sigma_{+}^{\infty}X)\in\operatorname{CAlg}(\operatorname{Sp}_{G})$. Since separability can be checked in the homotopy category, it follows from [BDS15, Theorem 1.1] that the commutative algebra $\mathbb{D}(\Sigma_{+}^{\infty}X)$ is separable in $\operatorname{Sp}_{G}$. Given $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$, there is a base change functor
$R\otimes-\colon\operatorname{Sp}_{G}\to\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)$
which is symmetric monoidal. We let $R^{X}$ denote the image of $\mathbb{D}(\Sigma_{+}^{\infty}X)$ under this functor, and observe that $R^{X}$ is a separable commutative algebra which is compact as an $R$-module. We therefore obtain a functor of groupoids
(9.1) $R^{(-)}\colon\mathcal{F}\mathrm{in}^{\operatorname{op}}_{G}\to\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega}).$
More generally, given a family $\mathscr{F}$ of subgroups of $G$, we have a functor
(9.2) $R^{(-)}_{/\mathscr{F}}\colon\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}/\mathscr{F}\subseteq\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}\stackrel{{\scriptstyle R^{(-)}}}{{\longrightarrow}}\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega})\longrightarrow\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}),$
$X\mapsto R\otimes\widetilde{E}\mathscr{F}\otimes\mathbb{D}(\Sigma_{+}^{\infty}X).$
Note that Equation 9.2 generalizes Equation 9.1 because $R^{(-)}_{/\varnothing}=R^{(-)}$.
Terminology 9.3.
We refer to the separable algebras in the essential image of (9.2) as standard.
Now assume that $K\subseteq G$ is a subgroup such that $K\not\in\mathscr{F}$ and $K^{\prime}\in\mathscr{F}$ for all $K^{\prime}\subsetneq K$. We note that the canonical inclusion $i_{K}\colon\mathcal{F}\mathrm{in}_{G}/\mathscr{F}\cup K\to\mathcal{F}\mathrm{in}_{G}/\mathscr{F}$ admits a retraction
$\pi_{K}\colon\mathcal{F}\mathrm{in}_{G}/\mathscr{F}\to\mathcal{F}\mathrm{in}_{G}/\mathscr{F}\cup K$
which, using the decomposition given in 7.8, sends the orbit $G/K$ to the empty set. The next result illustrates how the functors $R^{(-)}_{/\mathscr{F}}$ and $R^{(-)}_{/\mathscr{F}\cup K}$ relate to one another.
Lemma 9.4.
In the situation above, we have $R_{/\mathscr{F}\cup K}^{(-)}\simeq L_{\mathscr{F}\cup K}R_{/\mathscr{F}}^{i_{K}(-)}$ and the following diagram commutes
[…]
Proof.
The first claim follows by unravelling the definitions and using the equivalence $\widetilde{E}(\mathscr{F}\cup K)\xrightarrow{\sim}\widetilde{E}\mathscr{F}\otimes\widetilde{E}(\mathscr{F}\cup K)$. For the commutativity of the square we note that by the same argument as in Lemma 7.10, there is a factorization of $L_{\mathscr{F}\cup K}\circ R_{/\mathscr{F}}^{(-)}$ through $\pi_{K}^{\operatorname{op}}$, and by the first part of the lemma, this factorization must be through $R_{/\mathscr{F}\cup K}^{(-)}$.
As already discussed in the introduction, our goal in this paper is to find conditions to ensure that the functor (Equation 9.2) is an equivalence. The next result enables an induction on $\mathcal{F}$.
Theorem 9.5.
Let $G$ be a finite group, $\mathscr{F}$ a family of subgroups of $G$ and let $K\subseteq G$ be a subgroup such that $K\not\in\mathscr{F}$ and $K^{\prime}\in\mathscr{F}$ for all $K^{\prime}\subsetneq K$. Consider $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$ and assume that $\Phi^{K}(R)\in\operatorname{Fun}(W_{G}(K),\operatorname{CAlg})$ satisfies IC, RC and is separably closed. Assume furthermore that the functor
$R^{(-)}_{/\mathscr{F}\cup K}\colon(\mathcal{F}\mathrm{in}_{G}/\mathscr{F}\cup K)^{\operatorname{op}}\xrightarrow{\sim}\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathcal{F}\cup K)^{\omega})$
is an equivalence. Then the functor
$R^{(-)}_{/\mathscr{F}}\colon(\mathcal{F}\mathrm{in}_{G}/\mathscr{F})^{\operatorname{op}}\xrightarrow{\sim}\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathcal{F})^{\omega})$
is also an equivalence.
Proof.
For the duration of this proof we set $\Gamma:=W_{G}(K)$. The core of the proof is the following diagram which we will explain:
[…]
The right most face of the cube is obtained by applying $\mathcal{C}\mathcal{S}\mathrm{ep}(-)$ to the pullback square of symmetric monoidal stable $\infty$-categories of Theorem 8.2. In particular, this is a pullback square. Consider the left most face of the cube: this commutes because for $X\in\mathcal{F}\mathrm{in}_{G}$, the $K$-fixed points $X^{K}$ is a free $\Gamma$-set if $X\simeq\coprod_{i}G/K$ in $\mathcal{F}\mathrm{in}_{G}$. We claim that this left face of the cube is a pullback. This follows from the pasting lemma, [Lur09, Proposition 4.4.2.1], since the bottom left square and the outer square are pullback squares, for the latter use 7.8.
Next, we explain how the remaining faces of the cube arise and why they commute. The top and bottom faces are given by the squares in Lemma 9.4 and Lemma 7.10, so they commute. In particular we deduce that $\iota$ is fully faithful and that the functor $(\Phi^{K}R)^{(-)}$ is an equivalence, as depicted in the diagram. The front face of the cube commutes since
$\Phi^{K}R_{/\mathscr{F}}^{(X)}$ $=\Phi^{K}(R\otimes\widetilde{E}\mathscr{F}\otimes\mathbb{D}(\Sigma_{+}^{\infty}X))$
$\simeq\Phi^{K}(R)\otimes\mathbb{D}(\Sigma_{+}^{\infty}X^{K})$
$\simeq\underline{\mathrm{Hom}}(\Sigma_{+}^{\infty}X^{K},\Phi^{K}(R))$
$\simeq\prod_{X^{K}}\Phi^{K}(R)$
$\simeq(\Phi^{K}R)^{X^{K}}$
where we used the properties of the geometric fixed points, dualizability of $\Sigma_{+}^{\infty}X^{K}$, the properties of the cotensor functor. For the back face of the cube we use the first claim in Lemma 9.4 and the commutativity of the other faces to compute:
$\overline{\Phi}^{K}R^{X}_{/\mathscr{F}\cup K}$ $\simeq\overline{\Phi}^{K}L_{\mathscr{F}\cup K}R^{i_{K}(X)}_{/\mathscr{F}}$
$\simeq q\Phi^{K}R^{i_{K}(X)}_{\mathscr{F}}$
$\simeq q(\Phi^{K}R)^{i_{K}(X)^{K}}$
$\simeq\iota\operatorname{proj}(i_{K}(X)^{K})$
$\simeq\iota(\pi_{K}i_{K}(X))^{K}\simeq\iota(X)^{K}$
where in the last step we used fully faithfulness of $i_{K}$.
Now that we have shown that the above diagram commutes we are ready to prove that $R_{/\mathscr{F}}^{(-)}$ is an equivalence. Note that $R_{/\mathscr{F}}^{(-)}$ is fully faithful as it is a pullback of fully faithful functors. For essential surjectivity note that any object in $\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega})$ can be represented by a triple $(A,A^{\prime},f)$ where $A\in\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\Phi^{K}R)^{h\Gamma})$, $A^{\prime}\in\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}\cup K)^{\omega})$ and $f$ is an equivalence between the images of $A$ and $A^{\prime}$ in $\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\Phi^{K}R)^{t\Gamma})$. By our assumptions we can find preimages $X\in\mathcal{F}\mathrm{in}_{\Gamma}^{\mathrm{op}}$ and $X^{\prime}\in(\mathcal{F}\mathrm{in}_{G}/{\mathscr{F}\cup K})^{\mathrm{op}}$ of $A$ and $A^{\prime}$ respectively. Moreover using the commutativity of the back face of the cube and the fully faithfulness of $\iota$, we see that there is a unique equivalence $g$ between the images of $X$ and $X^{\prime}$, which hits $f$. Then the triple $(X,X^{\prime},g)$ defines an object of $(\mathcal{F}\mathrm{in}_{G}/\mathscr{F})^{\operatorname{op}}$ which hits $(A,A^{\prime},f)$ concluding the proof. ∎
Corollary 9.6.
Let $G$ be a finite group, $\mathscr{F}$ a family of subgroups of $G$ and consider $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$. Suppose that $\Phi^{K}(R)\in\operatorname{Fun}(W_{G}(K),\operatorname{CAlg})$ satisfies IC, RC and is separably closed for all subgroups $K\not\in\mathscr{F}$. Then the functor
$R_{/\mathscr{F}}^{(-)}\colon(\mathcal{F}\mathrm{in}_{G}/\mathscr{F})^{\operatorname{op}}\xrightarrow{\sim}\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F})^{\omega}).$
is an equivalence of $\infty$-groupoids.
Proof.
Firstly, we observe that the claim is trivial if $\mathscr{F}=\mathrm{All}$, so let us assume that $\mathscr{F}$ is a proper family. Choose an exhaustive filtration of families of subgroups of $G$
$\mathscr{F}=\mathscr{F}_{0}\subseteq\mathscr{F}_{1}\subseteq\ldots\subseteq\mathscr{F}_{i-1}\subseteq\mathscr{F}_{i}\subseteq\ldots\subseteq\mathscr{F}_{n}=\mathrm{All}$
with each $\mathscr{F}_{i}\backslash\mathscr{F}_{i-1}=\{(K_{i})\}$ consisting of a single conjugacy class. Note that this implies that for all proper subgroups $K\subsetneq K_{i}$ we have $K\in\mathscr{F}_{i-1}$. We will prove by descending induction on $i\leq n-1$, that the functor
$R_{/\mathscr{F}_{i}}^{(-)}\colon(\mathcal{F}\mathrm{in}_{G}/\mathscr{F}_{i})^{\operatorname{op}}\xrightarrow{\sim}\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}_{i})^{\omega})$
is an equivalence. The case $i=0$ will then be our claim. If $i=n-1$, then $\mathscr{F}_{i}=\mathscr{F}_{n-1}$ is the family of proper subgroups, so taking $G$-geometric fixed points induces an equivalence $\Phi^{G}\colon(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}_{n-1})^{\omega}\xrightarrow{\sim}\operatorname{Perf}(\Phi^{G}(R))$. Since $\Phi^{G}(R)$ is separably closed, we also deduce that
$\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathcal{F}_{n-1})^{\omega})=\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\Phi^{G}(R)))=\mathcal{F}\mathrm{in}^{\operatorname{op}}.$
By definition, any $X\in(\mathcal{F}\mathrm{in}_{G}/\mathscr{F}_{n-1})^{\operatorname{op}}$ is such that $X^{G}=X$, and so taking $G$-fixed points gives an equivalence $\mathcal{F}\mathrm{in}_{G}/\mathscr{F}_{n-1}\simeq\mathcal{F}\mathrm{in}$. It then follows that the composite
$(\mathcal{F}\mathrm{in}_{G}/\mathscr{F}_{n-1})^{\operatorname{op}}\xrightarrow{R_{/\mathscr{F}_{n-1}}^{(-)}}\mathcal{C}\mathcal{S}\mathrm{ep}((\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\mathscr{F}_{n-1})^{\omega})\xrightarrow[\sim]{\Phi^{G}}\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Perf}(\Phi^{G}R))\simeq\mathcal{F}\mathrm{in}^{\operatorname{op}}$
is an equivalence. By the $2$-out-of-$3$ property we conclude that $R_{/\mathscr{F}_{n-1}}^{(-)}$ is an equivalence proving the base case. The inductive step then follows from Theorem 9.5 and so we are done by induction. ∎
We record the two main examples we have. In the following we abbreviate
$\mathbb{Z}_{G}:=\operatorname{infl}_{e}^{G}\mathbb{Z}\in\operatorname{CAlg}(\operatorname{Sp}_{G})$
and recall that its category of modules can be identified with the $\infty$-category of derived Mackey functors of Kaledin, see [PSW22, Section 4].
Corollary 9.7.
Let $p$ be a prime number and $G$ a $p$-group. Then there are equivalences of $\infty$-groupoids
$\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}\xrightarrow{\sim}\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})\quad\mathrm{and}\quad\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}\xrightarrow{\sim}\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}).$
In particular, every separable commutative algebra in either $\operatorname{Sp}_{G}^{\omega}$ or $\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}$ is standard.
Proof.
If $G$ is trivial, the two claims follows from [NP24, Corollary 10.6]. If $G$ is a nontrivial $p$-group, then for all $K\subseteq G$, the $K$-geometric fixed points are $\Phi^{K}\mathbb{S}_{G}=\mathbb{S}$ and $\Phi^{K}\mathbb{Z}_{G}=\mathbb{Z}$ both with trivial Weyl group action. Note that these commutative algebras are all separably closed by [NP24, Corollary 10.6] again. They also satisfy IC and RC by Theorem 5.5, Lemma 6.5 and 5.4. So the claims follow from 9.6 applied to $\mathscr{F}=\varnothing$. ∎
We also describe the Galois group of these categories following [Mat16].
Corollary 9.8.
Let $p$ be a prime number and $G$ a $p$-group. Then the Galois groups of $\operatorname{Sp}_{G}^{\omega}$ and $\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}$ are trivial.
Proof.
To treat both cases simultaneously, let us write $A^{G}_{H}$ both for the separable commutative algebra $\mathbb{D}(G/H_{+})$ in $\operatorname{Sp}_{G}^{\omega}$ and for $\mathbb{Z}_{G}\otimes\mathbb{D}(G/H_{+})\in\operatorname{Mod}_{\operatorname{Sp}_{G}}(\mathbb{Z}_{G})^{\omega}$. First of all note that $\pi_{0}(\mathbb{1})$ in both categories is given by the Burnside ring $A(G)$, see [PSW22, Remark 2.20] for the second category. It follows that the two categories are indecomposable, see 10.1, and so their Balmer spectra are connected by [NP24, Proposition 3.7]. Since the finite Galois covers are exactly thoses separable algebras with locally constant (i.e. constant) degree function ([NP24, Corollary 8.13]), using 9.7 it will suffices to check that the degree function of $A^{G}_{H}$ is constant if and only if $G=H$. But in both cases the degree function is nonzero exactly at points in $\operatorname{supp}(A^{G}_{H})$ which is a proper subset of the Balmer spectrum unless $G=H$, see [PSW22, Lemma 2.12 and Proposition 2.13] and [BS17, Corollary 4.13]. ∎
10. Not all separable algebras are standard
In this section we show that not all separable commutative algebras in $\operatorname{Sp}_{G}^{\omega}$ are standard. We start off with the following:
Remark 10.1.
Consider $\mathbb{S}_{G}\in\operatorname{Sp}_{G}$, and note that $\pi_{0}^{G}(\mathbb{S}_{G})=A(G)$ the Burnside ring of the finite group $G$. By a result of Dress [Dre69], the ring $A(G)$ is indecomposable if and only if $G$ is solvable. It follows that if $G$ is not solvable then $\mathbb{S}_{G}$ decomposes non-trivially as a direct product $\mathbb{S}_{G}[f^{-1}]\times\mathbb{S}_{G}[(1-f)^{-1}]$ of separable algebras, which we claim are not standard. Indeed suppose that $\mathbb{S}_{G}[f^{-1}]\simeq\mathbb{D}(X_{+})$ for some finite $G$-set $X$. Since $\Phi^{e}\mathbb{S}_{G}=\mathbb{S}\in\operatorname{Sp}$ is indecomposable, we must have either $\Phi^{e}\mathbb{S}_{G}[f^{-1}]=0$ or $\Phi^{e}\mathbb{S}_{G}[(1-f)^{-1}]=0$. In the first case we deduce that $\Phi^{e}\mathbb{D}(\Sigma_{+}^{\infty}X)=0$ so $X=\varnothing$. This is a contradiction as we assumed that $\mathbb{S}_{G}[f^{-1}]\not=0$. In the second case we must have $\Phi^{e}\mathbb{S}_{G}[f^{-1}]=\mathbb{S}_{G}$. By checking on $e$-geometric fixed points we deduce that $X=\{\ast\}$, so that $\mathbb{S}_{G}[f^{-1}]\simeq\mathbb{S}_{G}$. This is a contradiction as $1-f$ is a nontrivial idempotent on $\mathbb{S}_{G}$ but trivial on $\mathbb{S}_{G}[f^{-1}]$.
However, even if the group is solvable (but not a $p$-group), there may exist non-standard separable algebras. To show this, we need some preliminary results.
Proposition 10.2.
Consider a pullback of groupoids
${\mathscr{A}}$${\mathscr{B}}$${\mathscr{C}}$${\mathscr{D}.}$$\scriptstyle{f}$$\scriptstyle{g}$
Then the canonical map $\varphi\colon\pi_{0}(\mathscr{A})\to\pi_{0}(\mathscr{B})\times_{\pi_{0}(\mathscr{D})}\pi_{0}(\mathscr{C})$ is surjective. Moreover given $a=(b,c,\eta\colon f(b)\xrightarrow{\sim}g(c))\in\pi_{0}(\mathscr{A})$, there is a bijection
[…]
where $\mathrm{Aut}_{\mathscr{B}}(b)$ acts on the right via $f$, and $\mathrm{Aut}_{\mathscr{C}}(c)^{\operatorname{op}}$ acts on the left via the group homomorphism
$\mathrm{Aut}_{\mathscr{C}}(c)^{\operatorname{op}}\to\mathrm{Aut}_{\mathscr{D}}(f(b)),\quad\gamma\mapsto\eta^{-1}g(\gamma)^{-1}\eta.$
Proof.
Pick $[b^{\prime},c^{\prime}]\in\pi_{0}(\mathscr{B})\times_{\pi_{0}(\mathscr{D})}\pi_{0}(\mathscr{C})$ so that $[f(b^{\prime})]=[g(c^{\prime})]$, that is, there exists an isomorphism $\eta\colon f(b^{\prime})\xrightarrow{\sim}g(c^{\prime})$ in $\mathscr{D}$. By definition of pullback in groupoids, the triple $(b^{\prime},c^{\prime},\eta^{\prime})$ defines an object $a^{\prime}\in\mathscr{A}$ which satisfies $\varphi(a^{\prime})=[b^{\prime},c^{\prime}]$, showing surjectivity of $\varphi$. For the second claim consider $a^{\prime}=[b^{\prime},c^{\prime},\eta^{\prime}\colon f(b^{\prime})\xrightarrow{\sim}g(c^{\prime})]\in\pi_{0}(\mathcal{A})$ and suppose that $\varphi(a)=\varphi(a^{\prime})$ which means that there exist isomorphisms $\beta\colon b\xrightarrow{\sim}b^{\prime}$ and $\gamma\colon c\xrightarrow{\sim}c^{\prime}$. We define a map (of sets)
[…]
$\quad a^{\prime}=[b^{\prime},c^{\prime},\eta^{\prime}]$ $\mapsto[\eta^{-1}g(\gamma)^{-1}\eta^{\prime}f(\beta)].$
We need to verify that this is well-defined so that the class in the double coset is independent of the chosen $\beta$ and $\gamma$. If we pick other isomorphisms $\beta^{\prime}\colon b\xrightarrow{\sim}b^{\prime}$ and $\gamma^{\prime}\colon c\xrightarrow{\sim}c^{\prime}$ witnessing that $\varphi(a)=\varphi(a^{\prime})$, then we can find $\nu\in\mathrm{Aut}_{\mathscr{B}}(b)$ and $\epsilon\in\mathrm{Aut}_{\mathscr{C}}(c)$ such that $\beta^{\prime}=\beta\nu$ and $\gamma^{\prime}=\gamma\epsilon$. Then the image under $\Phi$ is
$\eta^{-1}g(\gamma^{\prime})^{-1}\eta^{\prime}f(\beta^{\prime})$ $=\eta^{-1}g(\epsilon)^{-1}g(\gamma)^{-1}\eta^{\prime}f(\beta)f(\nu)$
$=\eta^{-1}g(\epsilon)^{-1}\eta\eta^{-1}g(\gamma)^{-1}\eta^{\prime}f(\beta)f(\nu)$
which defines the same class as $\eta^{-1}g(\gamma)^{-1}\eta^{\prime}f(\beta)$ in the codomain of $\Phi$. There is also a map
$\Psi\colon\mathrm{Aut}_{\mathscr{D}}(f(b))\to\varphi^{-1}\varphi(a),\quad\delta\mapsto[b,c,\eta\delta\colon f(b)\xrightarrow{\sim}g(c)].$
We first note that $\varphi\Psi(\delta)=[b,c]=\varphi(a)$ so $\Psi$ indeed lands in $\varphi^{-1}\varphi(a)$. To verify that $\Psi$ factors through the double coset we take $\delta\in\mathrm{Aut}_{\mathscr{D}}(f(b))$ and $\delta^{\prime}=\eta^{-1}g(\gamma)^{-1}\eta\delta f(\beta)$ for some $\beta\in\mathrm{Aut}_{\mathscr{B}}(b)$ and $\gamma\in\mathrm{Aut}_{\mathscr{C}}(c)$. We then need to show that $(b,c,\eta\delta)$ and $(b,c,\eta\delta^{\prime})=(b,c,g(\gamma)^{-1}\eta\delta f(\beta))$ are isomorphic in $\mathscr{A}$. This holds since there is a commutative diagram
[…]
hence the pair $(\beta,\gamma)$ defines the required isomorphism. Finally let us verify that $\Psi$ and $\Phi$ are inverse to one another. To this end, consider $a^{\prime}=[b^{\prime},c^{\prime},\eta^{\prime}]\in\varphi^{-1}\varphi(a)$ and choose isomorphisms $\beta\colon b\xrightarrow{\sim}b^{\prime}$ and $\gamma\colon c\xrightarrow{\sim}c^{\prime}$. Then we find that
$\Psi\Phi(a^{\prime})=[b,c,\eta\eta^{-1}g(\gamma)^{-1}\eta^{\prime}f(\beta)]=[b,c,g(\gamma)^{-1}\eta^{\prime}f(\beta)]$
which is isomorphic in $\mathscr{A}$ to $(b^{\prime},c^{\prime},\eta^{\prime})$ as there is a commutative square
[…]
showing that $\Psi\Phi(a^{\prime})=a^{\prime}$. For the other composite let $\delta\in\mathrm{Aut}_{\mathscr{D}}(f(b))$, then
$\Phi\Psi(\delta)=\Phi([b,c,\eta\delta])=[\eta^{-1}g(\gamma)^{-1}(\eta\delta)f(\beta)]=[\delta]$
since we can choose $\beta$ and $\gamma$ to be the corresponding identities. ∎
Lemma 10.3.
Let $\operatorname{\mathscr{C}}\in\operatorname{CAlg}(\mathrm{Pr}^{\mathrm{L}}_{\mathrm{st}})$ be such that $\pi_{0}(\mathbb{1})$ is indecomposable. For any $X\in\operatorname{Fin}$, we have
$\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(\mathbb{1}^{X},\mathbb{1}^{X})\simeq\operatorname{Map}_{\operatorname{Fin}}(X,X).$
In particular, the automorphism group of the commutative algebra $\mathbb{1}^{X}$ is given by the symmetric group $\Sigma_{X}$.
Proof.
Note that $\mathbb{1}^{X}=\prod_{X}\mathbb{1}$. Therefore using 5.1, we calculate
$\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(\mathbb{1}^{X},\mathbb{1}^{X})$ $=\coprod_{X}\prod_{X}\operatorname{Map}_{\operatorname{CAlg}(\operatorname{\mathscr{C}})}(\mathbb{1},\mathbb{1})$
$=\coprod_{X}\prod_{X}\operatorname{Map}_{\operatorname{Fin}}(\ast,\ast)$
$=\operatorname{Map}_{\operatorname{Fin}}(X,X).$
Example 10.4.
Let $G=C_{6}$ be the cyclic group of order $6$. We take the following exhaustive collection of families
$\varnothing\subseteq\{e\}\subseteq\{e,C_{2}\}\subseteq\{e,C_{2},C_{3}\}\subseteq\mathrm{All}.$
Using 8.3 with $\mathcal{F}=\varnothing$ we obtain the following pullback square of groupoids:
${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})}$${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega}/\{e\})}$${\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Fun}(BG,\operatorname{Sp}^{\omega}))}$${\mathcal{C}\mathcal{S}\mathrm{ep}(\mathrm{StMod}^{\mathrm{fin}}(\mathbb{S}_{2}^{\wedge}G))\times\mathcal{C}\mathcal{S}\mathrm{ep}(\mathrm{StMod}^{\mathrm{fin}}(\mathbb{S}_{3}^{\wedge}G)).}$$\scriptstyle{L^{\omega}_{\{e\}}}$$\scriptstyle{(\widetilde{\Phi}^{e})^{\omega}}$$\scriptstyle{\beta\circ(\overline{\Phi}^{e})^{\omega}}$$\scriptstyle{\beta\circ\pi}$
Using 9.6 and the fact that the corresponding Weyl groups are either $2$-groups or $3$-groups, we see that
$\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega}/\{e\})\simeq(\mathcal{F}\mathrm{in}_{G}/\{e\})^{\operatorname{op}},$
i.e., all separable algebras in $\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega}/\{e\})$ are standard. Using Lemma 7.6, we also find that
$\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Fun}(BG,\operatorname{Sp}^{\omega}))\simeq\mathcal{F}\mathrm{in}_{G}^{\operatorname{op}}.$
We also observe that $(\mathbb{S}_{2}^{\wedge})^{tG}$ and $(\mathbb{S}_{3}^{\wedge})^{tG}$ are indecomposable (argue as in the proof of Theorem 5.5), so the stable module categories appearing in the bottom right corner have indecomposable unit. By Proposition 10.2 any object $A\in\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})$ can be represented by a tuple
$A=(A_{1},A_{2},\eta_{1}\colon A_{1}^{\prime}\simeq A_{2}^{\prime},\eta_{2}\colon A_{1}^{\prime\prime}\simeq A_{2}^{\prime\prime})$
with
  • (1)
    an object $A_{1}\in\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega}/\{e\})$;
  • (2)
    an object $A_{2}\in\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Fun}(BG,\operatorname{Sp}^{\omega}))$;
  • (3)
    an isomorphism $\eta_{1}\colon A_{1}^{\prime}\simeq A_{2}^{\prime}$ between the images of $A_{1}$ and $A_{2}$ in $\mathcal{C}\mathcal{S}\mathrm{ep}(\mathrm{StMod}^{\mathrm{fin}}(\mathbb{S}_{2}^{\wedge}G))$;
  • (4)
    an isomorphism $\eta_{2}\colon A_{1}^{\prime\prime}\simeq A_{2}^{\prime\prime}$ between the images of $A_{1}$ and $A_{2}$ in $\mathcal{C}\mathcal{S}\mathrm{ep}(\mathrm{StMod}^{\mathrm{fin}}(\mathbb{S}_{3}^{\wedge}G))$.
We now claim that $A\in\mathcal{C}\mathcal{S}\mathrm{ep}(\operatorname{Sp}_{G}^{\omega})$ determined by $A_{1}=\mathbb{1}^{\times 2}$, $A_{2}=\mathbb{1}^{\times 2}$, $\eta_{1}=\mathrm{id}$ and $\eta_{2}=\sigma\in\Sigma_{2}$ the flip isomorphism (see Lemma 10.3) is not standard: Indeed, since $(\tilde{\Phi}^{e})^{\omega}$ determines the $G$-set defining a standard algebra, we would have to have $A\simeq\mathbb{1}^{\times 2}$, which by Proposition 10.2 would imply
[…]
which is not true (observe both maps of $\Sigma_{2}$ are the diagonal embedding).
Remark 10.5.
The argument in the previous example also applies to $\mathbb{Z}_{G}=\operatorname{infl}_{e}^{G}\mathbb{Z}$. It follows that in the category of compact derived $C_{6}$-Mackey functor not all separable commutative algebras are standard.
11. Other examples
In this final section we record some other classification results.
11.1. Free $G$-spectra
Consider $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$ such that $R\otimes\widetilde{E}G\simeq 0$ and $\Phi^{e}(R)$ separably closed. An example to keep in mind is Atiyah’s K-theory with reality where $G=C_{2}$, see [BCH+23, Example 8.15]; note that $\Phi^{e}KR=KU$ is separably closed by [NP24, Theorem 12.9]. Consider the pullback square from Theorem 8.2 with $R$ as above, $\mathcal{F}=\varnothing$ and $K=e$. The right top vertex of the square is
$(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)/\{e\})^{\omega}\simeq\operatorname{Mod}_{\operatorname{Sp}_{G}}(R\otimes\widetilde{E}G)^{\omega}\simeq 0$
It follows that the functor $\Phi^{e}\colon\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega}\to\operatorname{Perf}(\Phi^{e}R)^{hG}$ is fully faithful. Passing to separable algebras we obtain a fully faithful functor
$\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega})\hookrightarrow\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(\Phi^{e}(R)))^{hG}\simeq\operatorname{Fin}_{G}^{\operatorname{op}}$
by our assumption on $\Phi^{e}(R)$. But $R\otimes\mathbb{D}(\Sigma_{+}^{\infty}X)\in\operatorname{CAlg}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(R)^{\omega})$ for $X\in\operatorname{Fin}_{G}$ are always separable, so the above map is an equivalence.
11.2. Borel $G$-spectra
Consider $R\in\operatorname{CAlg}$ and its associated Borel $G$-spectrum $\underline{R}\in\operatorname{CAlg}(\operatorname{Sp}_{G})$. Then it follows from [MNN17, Corollary 6.21] that there is a fully faithful functor
$\operatorname{Mod}_{\operatorname{Sp}_{G}}(\underline{R})^{\omega}\hookrightarrow\operatorname{Fun}(BG,\operatorname{Mod}(R))^{\omega}\subseteq\operatorname{Fun}(BG,\operatorname{Perf}(R)).$
given by $\Phi^{e}$. Passing to separable algebras we find that
$\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Mod}_{\operatorname{Sp}_{G}}(\underline{R})^{\omega})\hookrightarrow\operatorname{Fun}(BG,\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(R))).$
If $R$ is separably closed, so that $\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(R))\simeq\operatorname{Fin}^{\operatorname{op}}$, then the above functor is also essentially surjective since $\underline{R}\otimes\mathbb{D}(\Sigma_{+}^{\infty}X)$ is a separable commutative algebra in $\operatorname{Mod}_{\operatorname{Sp}_{G}}(\underline{R})^{\omega}$ hitting the finite finite $G$-set $X$.